OpenStudy (anonymous):

Rationalinzing Expressions: x^2 +2x over 12x +54 - 3-x over 8x +36

OpenStudy (anonymous):

x^2 + 2x = x(x+2) 1 _______________ x ____________________ 11x + 51 4(2x + 9) Then simplify

OpenStudy (campbell_st):

is the problem \[\frac{ x^2 + 12x}{12x - 54} - \frac{3 - x}{8x + 36}\]

OpenStudy (anonymous):

yes @campbell_st

OpenStudy (anonymous):

Just kidding I meant x^2 + 2x

OpenStudy (anonymous):

it should look like \[x^2 + 2x \over 12x + 54 \]

OpenStudy (campbell_st):

thank you.... then with some factorizing you get \[\frac{x(x+12)}{6(2x - 9)} - \frac{3 - x}{4(2x + 9)}\] multiply the denoninators to get a common denominator and cross multiply the numerators \[\frac{x \times(x+12) \times 4\times(2x+9) - (3 - x) \times 6\times(2x-9)}{6(2x-9)\times 4(2x +9)}\] you'll need to expand and collect like terms... the denominator is the difference of 2 squares. \[\frac{4x \times (2x^2 + 33x + 81) - 6\times (-2x^2 +15x - 27)}{24\times(4x^2 - 81)}\] simplify and collect like terms \[\frac{ 8x^3 + 144x^2 +234x + 162}{24\times(4x^2 - 81)}\] \[\frac{4x^3 + 72x^2 + 117x + 81}{12(4x^2 - 81)}\]

OpenStudy (anonymous):

@campbell_st , the original number was 2 not 12. It would completley change the equation wouldnt it?

OpenStudy (campbell_st):

yes it would